OpenStudy (anonymous):

A manufacturer of shipping boxes has a box shaped like a cube. The side length is (5a + 4b) . What is the volume of the box in terms of a and b?

OpenStudy (calculusxy):

\[\large (5a + 4b)^3\] it is like: \[\large (5a+4b)(5a + 4b)(5a + 4b)\] i would use FOIL so let's start with \[\large (5a+4b)(5a+4b)\]

OpenStudy (calculusxy):

\[\large (5a \times 5a) + (5a \times 4b) + (4b \times 5a) + (4b \times 4b)\]

OpenStudy (calculusxy):

\[\large (5a + 4b)(25a^2 + 20ab + 20ba +16b^2)\]

OpenStudy (calculusxy):

Continue using FOIL

OpenStudy (anonymous):

whenever i watched the example video it had something to do with some kind of triangle not foil let me up load the video real quick

OpenStudy (calculusxy):

i think i understand this: